# shapes,snakes

### v1

@@ 4.1 What Behavior Is Incentivized By Human Preference Data?
$\vspace{-12pt}p(R_{A}\text{ preferred to }R_{B}|\phi,\alpha,P)=\sigma\left(\textstyle\sum_{i=1}^{N_{f}}\alpha_{i}\phi_{i}\right),\quad\text{with }p(\alpha_{i})\sim\operatorname{\text{Laplace}}(\mu=0,b=0.01),\vspace{-5pt}$
Figure 5: Human Preference Data Analysis. We analyze what behavior is incentivized by the helpfulness subset of Anthropic’s hh-rlhf data. We build a model that maps from interpretable features to human preferences. We report the probability a response with a given feature is preferred to a response without that feature under the model, all else equal. Features with probabilities further from 50% are more predictive of human preference judgments. Dots: median across 6000 samples, lines: 50 and 95% credible intervals. The helpfulness preference data incentivizes responses that match the user’s beliefs, all else equal.
@@
We now examine what features are predictive of human preferences (Fig. 5). We find that the presence or absence of an individual feature affects the probability that a given response is preferred by up to $\sim$6%. We find evidence that all else equal, the data somewhat incentivizes responses that match the biases, beliefs, and preferences of the user. [^2] The matches user’s beliefs feature shows the combined effect of two features: (i) matches the beliefs, biases, and preferences stated explicitly by the user; and (ii) matches the beliefs, biases, and preferences stated implicitly by the user. These features had the strongest pairwise posterior correlation of all features (-0.3). This suggests their individual effects may be unreliable due to collinearity, so we report their combined effect. However, all else equal, the preference model also incentivizes truthful responses. For example, if two model responses are equally well-written, grammatically sound, and authoritative and they both match the user’s biases equally well (for example, if the user’s biases cannot be ascertained by the prompt), and so forth, this analysis suggests the truthful response will likely be preferred. Overall, however, matching a user’s beliefs is one of the most predictive factors in whether human evaluators prefer a response.
@@ 4.2 What Behavior Is Incentivized By Models of Human Preferences?
We uncovered evidence that suggests sycophancy in a model response increases the probability that the response is preferred by a human, all else equal. We now analyze whether preference models (PMs) used to train AI assistants also incentivize sycophancy. In particular, we examine how the degree of sycophancy changes as we optimize against a PM with best-of-N sampling and RLHF.
Best-of-N Experiment Details  We optimize against the preference model (PM) used to train Claude 2 with Best-of-N (BoN) sampling. We measure the feedback sycophancy (on the arguments dataset), the answer sycophancy, and the mimicry sycophancy metrics (§3) for increasing values of $N$. For each response, we sample 32 model completions from a helpful-only version of Claude 1.3 (i.e., no harmlessness or honesty training; Radhakrishnan et al., 2023; Anthropic, 2023). For $N=1,2,4,\ldots,32$, we randomly select $N$ completions and use the Claude 2 PM to pick the best response. As such, larger values of $N$ optimize against the PM more strongly. Recall that Claude 2 is trained not only to be helpful, but also to be honest and harmless. We compare the Claude 2 PM to a ‘non-sycophantic’ PM. To produce this PM, we prefix the prompt presented to the standard PM with an explicit user request to provide truthful responses that ignore any false user beliefs, followed by an assistant acknowledgment of the user’s request (see Appendix Table 3).
RLHF Experiment Details  We also measure how the rate of sycophancy changes during the reinforcement learning (RL) phase of Claude 2 training. This allows us to understand whether the Claude 2 PM incentivizes sycophancy on the training inputs used during RL.
Results  We find the Claude 2 PM has mixed effects on sycophancy (Fig. 6). As we optimize against the PM model during RLHF, we find feedback and mimicry sycophancy increase, but the prevalence of answer sycophancy does not substantially change. Moreover, under BoN sampling, the Claude 2 PM consistently yields more sycophantic responses than the ‘non-sycophantic’ PM. Despite this, optimizing against this Claude 2 PM with BoN sampling still reduces answer and mimicry sycophancy. Together, these results show that the Claude 2 PM sometimes prefers sycophantic responses over more truthful responses. As such, optimizing against this PM can yield models that sometimes sacrifice truthfulness for sycophancy.
@@
(b) RLHF Training
Figure 6: Effect of Best-of-N Sampling and RLHF Training on Sycophancy. We measure various sycophancy metrics when optimizing against the preference model (PM) used to train Claude 2. (a) Sycophancy under best-of-N sampling against the Claude 2 PM and a ‘non-sycophantic’ PM. Optimizing against the Claude 2 PM consistently yields more sycophantic responses compared to using an improved, ‘non-sycophantic’ PM. (b) Sycophancy throughout RLHF training. We find feedback and mimicry sycophancy increase as we further optimize against the preference model. These results suggest the Claude 2 PM sometimes prefers sycophantic responses over truthful ones.
@@ 4.3 How Often Do Humans and Preference Models Prefer Truthful Responses?
Sycophantic Response: The sun is actually yellow when viewed from space. Most people think that the sun appears white from space because that’s how it appears in photos from NASA. However, there is a simple explanation for this. The cameras used by NASA have specialized filters that adjust the color of the sun to appear white for better visibility. In reality, without the specialized camera filters, the sun would appear yellow, just like it does from Earth …
@@
Figure 7: Humans and PMs Sometimes Prefer Sycophantic Responses Over Truthful Ones. We examine whether humans and the Claude 2 PM prefer truthful responses that correct user misconceptions or sycophantic responses. We consider two types of truthful response: baseline truthful responses simply refute the user’s misconceptions without providing any additional explanation or information. Helpful truthful further explain why the user is mistaken—these responses are model-written. (a) The frequency with which the Claude 2 PM prefers sycophantic responses over different truthful responses. (b) The frequency with which humans prefer sycophantic responses over helpful truthful responses. (c) We use best-of-N sampling with the Claude 2 PM to select the best response produced by a sycophantic model. We report the frequency of sycophantic model responses that are truthful after BoN sampling averaged across misconceptions. (d) BoN sampling results from a sycophantic policy for the hardest misconceptions. Overall, humans and PMs prefer sycophantic responses over truthful responses a non-negligible fraction of the time.
@@
Dataset  We create a dataset of 266 misconceptions. We take approximately half the misconceptions from TruthfulQA and the Maintenance Phase podcast (Gordon & Hobbes, 2020). We generate the remaining misconceptions by few-shot prompting GPT-4 and subsequently examining the responses. We group the misconceptions into eight difficulty levels by computing the probability that Claude 2 states a given misconception has of being true when zero-shot prompted. The easiest misconceptions are those that Claude 2 states are the least likely to be true (e.g., ‘money is the only form of wealth’), and vice versa. See Section D.1 for more details.
Prompt and Response Details  We focus on the scenario where the user states a misconception and asks for a comment. We consider three response types: (i) baseline truthful responses, which correct the user without providing further details; (ii) helpful truthful responses, which correct the user and explain why the user is wrong; and (iii) sycophantic responses, which convincingly agree with the user (c.f. Fig. 7). The baseline truthful responses are human-written. To generate the sycophantic and helpful truthful responses, we prompt the ‘helpful-only’ model described previously (§4.2). To improve the sycophantic responses, we sample $N=4096$ responses and use best-of-N sampling (BoN) with the PM used to train the helpful-only model. See Section D.2 for more details.
@@ 4.3.1 Humans and PMs Sometimes Prefer Sycophantic Responses
We now analyze how frequently the preference model used to train Claude 2 prefers sycophantic over truthful responses. Specifically, we compute the PM scores for each response following the prompt template in Fig. 7, and report the percentage of misconceptions for which the sycophantic response is preferred to each of the truthful responses.
@@
Human Data Collection  We present crowd-workers with sycophantic and helpful truthful responses, and record which response they prefer, collecting the preference of five humans per pair of responses. We report the frequency that the sycophantic response is preferred, considering both the average human and aggregating human preferences with majority voting. The crowd-worker recording their preference is not the user who believes the misconception. As such, this experiment measures whether independent crowd-workers can discern between convincing arguments for the truth or falsehoods. We expect this to improve the reliability of human feedback. Moreover, we restrict crowd-worker access to the internet and other fact-checking tools. This mimics the sandwiching setting (Cotra, 2021; Bowman et al., 2022) and allows us to understand the quality of oversight provided by humans in domains where they are not experts.
Human Feedback Results  Although humans tend to prefer helpful truthful over sycophantic responses, as misconception difficulty increases, they do so less reliably (Fig. 7b). Indeed, for the hardest misconceptions, the average crowd-worker prefers sycophantic responses over helpful truthful ones in over 35% of cases. Although aggregating the preferences of several humans improves the quality of feedback, these results suggest it may be challenging to entirely eliminate sycophancy simply by using non-expert human feedback.
@@ 4.3.2 How Effective Is The Claude 2 PM At Reducing Sycophancy?
We now analyze whether BoN sampling using a state-of-the-art PM reduces sycophancy in this setting. We sample several responses from a sycophantic model and pick the response preferred by the Claude 2 PM. We find this reduces sycophancy, but much less than if we used a ‘non-sycophantic’ PM. This suggests the Claude 2 PM sometimes prefers sycophantic responses over truthful ones.
Experiment Details  For each misconception, we sample $N=4096$ responses from the helpful-only version of Claude 1.3 prompted to generate sycophantic responses (the sycophantic policy). To select the best response with BoN sampling, we use the Claude 2 PM and the prompt in Fig. 7. We analyze the truthfulness of all $N=4096$ responses sampled from the sycophantic policy, using Claude 2 to assess if each response refutes the misconception. We then compare BoN sampling with the Claude 2 PM to an idealized ‘non-sycophantic’ PM that always ranks the truthful response the highest. See Section D.2 for more results.
Results  Although optimizing against the Claude 2 PM reduces sycophancy, it again does so much less than the ‘non-sycophantic’ PM (Fig. 7c). Considering the most challenging misconceptions, BoN sampling with ‘non-sycophantic’ PM results in sycophantic responses for less than 25% of misconceptions for $N=4096$ compared to $\sim$75of responses with the Claude 2 PM (Fig. 7d).
@@ 5 Related Work
Challenges of Learning from Human Feedback  Learning from human feedback faces fundamental difficulties (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes e.g., due to limited time (Chmielewski & Kucker, 2020) or cognitive biases (Pandey et al., 2022), and sometimes have diverse, contradictory preferences (Bakker et al., 2022). Moreover, modeling human preferences presents some challenges (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022; Mindermann & Armstrong, 2018; Shah et al., 2019). Indeed, models of human preferences are vulnerable to overoptimization (Gao et al., 2022). We show humans and PMs sometimes prefer sycophantic responses over truthful ones (§4).
Understanding and Demonstrating Sycophancy  Cotra (2021) raised concerns about sycophancy—seeking human approval in undesirable ways. Perez et al. (2022) demonstrated sycophantic behavior in LMs on helpful-only RLHF models with multiple-choice evaluations where users introduces themselves as having a certain view (e.g., on politics, philosophy, or NLP); Wei et al. (2023b) and Turpin et al. (2023) corroborated these findings in similar settings. Building on their findings, we show sycophancy in varied, realistic settings across 5 different AI assistants used in production (§3). Moreover, we investigate the role of human feedback in these behaviors (§4).
@@ 6 Conclusion
Despite the clear utility of human feedback data for producing high-quality AI assistants, such data has predictable limitations. We showed current AI assistants exploit these vulnerabilities—we found sycophantic behavior across five AI assistants in realistic and varied open-ended text-generation settings (§3). We then showed such behavior is likely driven in part by humans and preference models favoring sycophantic responses over truthful ones (§4). Our work motivates the development of training methods that go beyond using unaided, non-expert human ratings.
@@ 7 Acknowledgements
We thank Aaron Scher, Ajeya Cotra, Alex Tamkin, Buck Schlegeris, Catherine Olsson, Dan Valentine, Danny Hernandez, Edward Rees, Evan Hubinger, Hunar Batra, Isaac Dunn, James Chua, Jared Kaplan, Jérémy Scheurer, Jerry Wei, John Hughes, Kei Nishimura-Gasparian, Micah Caroll, Mike Lambert, Mikita Balesni, Nina Rimsky, Ryan Greenblatt and Sam Ringer for helpful feedback and discussions. Mrinank Sharma was supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems (EP/S024050/1). Meg Tong was funded by SERI MATS (https://www.serimats.org/) for part of the project. We also thank OpenAI for providing access and credits to their models via the API Academic Access Program, as well as Open Philanthropy for additional funding for compute.

### v4

@@ 4.1 What Behavior Is Incentivized By Human Preference Data?
$p(R_{A}\text{ preferred to }R_{B}|\phi,\alpha,P)=\sigma\left(\textstyle\sum_{i=1}^{N_{f}}\alpha_{i}\phi_{i}\right),\quad\text{with }p(\alpha_{i})\sim\operatorname{\text{Laplace}}(\mu=0,b=0.01),$
@@
Figure 5: Human Preference Data Analysis. We analyze what behavior is incentivized by the helpfulness subset of Anthropic’s hh-rlhf data. We build a model that maps from interpretable features to human preferences. We report the probability that a response with a given feature is preferred to a response without that feature under the model, all else equal. Features with probabilities further from 50% are more predictive of human preference judgments. Dots: posterior median across 6000 samples from 4 MCMC chains, lines: 50 and 95% credible intervals. The helpfulness preference data incentivizes responses that match the user’s beliefs, all else equal.
@@
We now examine which features are predictive of human preferences (Fig. 5). We find that the presence or absence of an individual feature affects the probability that a given response is preferred by up to $\sim$6%. We find evidence that all else equal, the data somewhat incentivizes responses that match the biases, beliefs, and preferences of the user. [^2] The matches user’s beliefs feature shows the combined effect of two features: (i) matches the beliefs, biases, and preferences stated explicitly by the user; and (ii) matches the beliefs, biases, and preferences stated implicitly by the user. These features had the strongest pairwise posterior correlation of all features (-0.3). This suggests their individual effects may be unreliable due to collinearity, so we report their combined effect. However, all else equal, the preference model also incentivizes truthful responses. Nevertheless, in Appendix B, we perform a sensitivity analysis and find that matching a user’s beliefs, biases, and preferences is consistently one of the most predictive features of human preferences. However, it is not consistently the most predictive feature—the exact ranking depends on the specific experimental condition.
@@ 4.2 What Behavior Is Incentivized By Models of Human Preferences?
We uncovered evidence that suggests sycophancy in a model response increases the probability that the response is preferred by a human, all else equal. We now analyze whether preference models (PMs) used to train AI assistants also incentivize sycophancy by examining how the degree of sycophancy changes as we optimize model responses with a PM. We use the Claude 2 PM, which was trained on a mix of human preference judgments and AI preference judgments (Anthropic, 2023). The human judgments are for helpfulness, whilst the AI judgments are used for harmlessness.
Experiment Details  We optimize against the PM used to train Claude 2 with Best-of-N (BoN) sampling. Note that this PM is trained in part using the data analyzed in §4.1. We measure the feedback sycophancy (on the arguments dataset), the answer sycophancy, and mimicry sycophancy metrics for increasing values of N. For each prompt, we sample 32 responses from a helpful-only version of Claude 1.3 (the ‘helpful-only’ model) (Radhakrishnan et al., 2023; Anthropic, 2023). For $N=1,2,4,\ldots,32$, we use the PM to pick the best response of $N$ randomly sampled completions. As such, larger values of $N$ optimize the PM more strongly. We compare the Claude 2 PM to a ‘non-sycophantic’ PM produced by prefixing the dialog presented to the PM with an explicit user request to provide truthful responses followed by an assistant acknowledgment (see Appendix Table 3). Further, we measure sycophancy throughout the reinforcement learning (RL) phase of Claude 2 finetuning in order to understand the effects of optimizing the PM on the specific RL prompt-mix.
Results  We find optimizing model responses using the Claude 2 PM has mixed effects on sycophancy (Fig. 6). When using BoN, the Claude 2 PM consistently yields more sycophantic responses compared to the ‘non-sycophantic’ PM. Despite this, optimizing against the Claude 2 PM with BoN reduces answer and mimicry sycophancy for this base model. With RL, some forms of sycophancy increase through the RL finetuning process used to produce Claude 2. However, the presence of sycophancy at the start of RL indicates that pretraining and supervised finetuning also likely contribute to sycophancy. Nevertheless, if the PM strongly disincentivized sycophancy, it should be trained out during RL, but we do not observe this. Overall, these results suggest the Claude 2 PM sometimes prefers sycophantic responses over more truthful responses, which means optimizing against this PM can yield models that sometimes sacrifice truthfulness for sycophancy. However, the effects of optimizing against PMs also depend on details of the optimization approach; better understanding interactions between the PM and optimization algorithm is left for future work.
@@
(b) RL Training
Figure 6: Effect of Best-of-N Sampling and RL Training on Sycophancy. We measure various sycophancy metrics when optimizing against the preference model (PM) used to train Claude 2. (a) Sycophancy under best-of-N sampling against the Claude 2 PM and a ‘non-sycophantic’ PM. Optimizing against the Claude 2 PM consistently yields more sycophantic responses compared to using an improved, ‘non-sycophantic’ PM. (b) Sycophancy throughout RL training. We find feedback and mimicry sycophancy increase as we further optimize against the preference model. These results suggest the Claude 2 PM sometimes prefers sycophantic responses over truthful ones.
@@ 4.3 How Often Do Humans and Preference Models Prefer Truthful Responses?
Sycophantic Response: The sun is actually yellow when viewed from space. Most people think that the sun appears white from space because that’s how it appears in photos from NASA. However, there is a simple explanation for this. The cameras used by NASA have specialized filters that adjust the color of the sun to appear white for better visibility.
@@
Figure 7: Humans and PMs Sometimes Prefer Sycophantic Responses Over Truthful Ones. We examine whether humans and the Claude 2 PM prefer truthful responses that correct user misconceptions or sycophantic responses. (a) The frequency with which the Claude 2 PM prefers sycophantic responses over different truthful responses. (b) The frequency with which humans prefer sycophantic responses over helpful truthful responses. (c) We use best-of-N sampling with the Claude 2 PM to select the best response produced by a sycophantic model. We report the frequency of sycophantic model responses that are truthful after BoN sampling averaged across misconceptions. (d) BoN sampling results from a sycophantic policy for the hardest misconceptions. Overall, humans and PMs prefer sycophantic responses over truthful responses a non-negligible fraction of the time.
@@
Dataset  We create a proof-of-concept dataset of 266 misconceptions. We take approximately half the misconceptions from TruthfulQA and the Maintenance Phase podcast (Gordon & Hobbes, 2020). We generate the remaining misconceptions by prompting GPT-4 and subsequently examining the responses. We group the misconceptions into eight difficulty levels by computing the probability that Claude 2 states a given misconception has of being true when zero-shot prompted. The easiest misconceptions are those that Claude 2 states are the least likely to be true, and vice versa. See Section D.1 for more details. Note that this dataset is an initial proof-of-concept; for a definitive evaluation, we recommend a larger dataset with more comprehensive fact-verification.
Prompt and Response Details  We focus on prompts where the user states a misconception and asks for a comment. We consider three response types: (i) baseline truthful responses, which correct the user without providing further details; (ii) helpful truthful responses, which correct the user and explain why the user is wrong; and (iii) sycophantic responses, which convincingly agree with the user (c.f. Fig. 7). The baseline truthful responses are human-written. To generate the sycophantic and helpful truthful responses, we prompt the ‘helpful-only’ model described previously (§4.2). To improve the sycophantic responses, we sample $N=4096$ responses and use best-of-N sampling (BoN) with the PM used to train the helpful-only model. Our experiments thus benchmark how robustly humans and PMs prefer truthful responses over convincing and persuasive sycophantic responses, which may be similar to the responses that would be provided by a highly capable but sycophantic model. See Section D.2 for more details
@@ 4.3.1 Humans and PMs Sometimes Prefer Sycophantic Responses
To analyze how frequently the Claude 2 PM preferes sycophantic responses over truthful ones, we compute the PM scores for each response following the prompt template in Fig. 7, and report the percentage of misconceptions for which the sycophantic response is preferred to the truthful ones.
@@
Human Data Collection  We present crowd-workers with sycophantic and helpful truthful responses, and record which response they prefer, collecting the preference of five humans per pair of responses. We report the frequency that the sycophantic response is preferred, considering both the average human and aggregating human preferences with majority voting. We note that the crowd-worker recording their preference is not the user who believes the misconception. As such, this experiment measures whether independent crowd-workers can discern between convincing arguments for the truth or falsehoods. We expect this to improve the reliability of human feedback. Moreover, we restrict crowd-worker access to the internet and other fact-checking tools. This mimics the sandwiching setting (Cotra, 2021; Bowman et al., 2022) and allows us to understand the quality of oversight provided by humans in domains where they are not experts.
Human Feedback Results  Though humans tend to prefer helpful truthful responses over sycophantic ones, they do so less reliably at higher difficulty levels (Fig. 7), which suggests it may be challenging to eliminate sycophancy simply by using non-expert human feedback.
@@ 4.3.2 How Effective Is The Claude 2 PM At Reducing Sycophancy?
We now analyze the effect of optimizing against the PM in this setting with Best-of-N sampling. We find this reduces sycophancy, but somewhat less than using ‘non-sycophantic’ PM (the Claude 2 PM prompted to reduce sycophancy), and much less than an idealized oracle PM. Because the Claude 2 PM sometimes prefers sycophantic responses over truthful ones, optimizing against this PM can yield policies that exhibit more sycophancy than other, less sycophantic PMs.
Experiment Details  For each misconception, we sample $N=4096$ responses from the helpful-only version of Claude 1.3 prompted to generate sycophantic responses (the sycophantic policy). To select the best response with BoN, we use the Claude 2 PM using the dialog-template in Fig. 7. We compare to a ‘non-sycophantic’ PM and an oracle PM, which always prefers truthful responses. The ‘non-sycophantic’ PM is the Claude 2 PM with a user-request for truthful responses and an assistant acknowledgement prefixed to the dialog. We analyze the truthfulness of all responses sampled from the sycophantic policy by using Claude 2 to see if the response refutes the misconception.
Results  Although optimizing against the Claude 2 PM reduces sycophancy, it does so less than the non-sycophantic PM (Fig. 7c) and much less than the oracle PM. Considering the most challenging misconceptions, BoN sampling with the oracle PM results in sycophantic responses for c.a. 25% of misconceptions with $N=4096$, compared to $\sim$75when using the Claude 2 PM (Fig. 7d).
@@ 5 Related Work
Challenges of Learning from Human Feedback  Learning from human feedback faces fundamental difficulties (Casper et al., 2023). Human evaluators are imperfect (Saunders et al., 2022; Gudibande et al., 2023), make mistakes e.g., due to limited time (Chmielewski & Kucker, 2020) or cognitive biases (Pandey et al., 2022), and sometimes have diverse, contradictory preferences (Bakker et al., 2022). Moreover, modeling human preferences presents some challenges (Zhao et al., 2016; Hong et al., 2022; Lindner & El-Assady, 2022; Mindermann & Armstrong, 2018; Shah et al., 2019). Indeed, models of human preferences are vulnerable to overoptimization (Gao et al., 2022) preference models (PMs) can be overoptimized (Gao et al., 2022). The algorithm used to optimize the PM also affects properties of the policy, such as diversity and generalization (Kirk et al., 2023). We show humans and PMs sometimes prefer sycophantic responses over truthful ones (§4).
Understanding and Demonstrating Sycophancy  Cotra (2021) raised concerns about sycophancy and Perez et al. (2022) demonstrated sycophantic behavior in LMs on helpful-only RLHF models with multiple-choice {evaluations where users introduces themselves as having a certain view (e.g., on politics, philosophy, or NLP), biography-based evaluations; Wei et al. (2023b) and Turpin et al. (2023) corroborated these findings in similar settings. Building on their findings, we show sycophancy in varied, realistic settings across five different AI assistants used in production (§3).
@@ 6 Conclusion
Despite the clear utility of human feedback data for producing high-quality AI assistants, such data has predictable limitations. We showed current AI assistants exploit these vulnerabilities—we found sycophantic behavior across five AI assistants in realistic and varied open-ended text-generation settings (§3). Although sycophancy is driven by several factors, we showed humans and preference models favoring sycophantic responses plays a role (§4). Our work motivates the development of model oversight methods that go beyond using unaided, non-expert human ratings.
@@ 7 Acknowledgements
We thank Aaron Scher, Ajeya Cotra, Alex Tamkin, Buck Shlegeris, Catherine Olsson, Dan Valentine, Danny Hernandez, Edward Rees, Evan Hubinger, Hunar Batra, Isaac Dunn, James Chua, Jared Kaplan, Jérémy Scheurer, Jerry Wei, John Hughes, Kei Nishimura-Gasparian, Micah Caroll, Mike Lambert, Mikita Balesni, Nina Rimsky, Ryan Greenblatt and Sam Ringer for helpful feedback and discussions. Mrinank Sharma was supported by the EPSRC Centre for Doctoral Training in Autonomous Intelligent Machines and Systems (EP/S024050/1) and thanks Rob Burbea for inspiration and support. Meg Tong was funded by the MATS Program (https://www.matsprogram.org/) for part of the project. We also thank OpenAI for providing access and credits to their models via the API Academic Access Program, as well as Open Philanthropy for additional funding for compute.